# Multi-task Photonic Reservoir Computing: Wavelength Division Multiplexing for Parallel Computing with a Silicon Microring Resonator

### v1

@@ 6.7.2 Impact of $\Delta\phi$ when addressing multiple, different tasks
Table 3: Comparison of the best performance of each task for different values of $\Delta\phi$ for a $\overline{P}_{\textrm{T}}$ range of $[-20,25]$ dBm and a $\Delta\omega/2\pi$ range of $[-100,100]$.
When addressing different tasks, changing the value of $\Delta\phi$ does not always have an impact on the performance behavior of each task over the parameter space as shown in Fig. 11 for the NARMA-10 task. In Fig. 11 (a,b,d), there is not much variation in the performance over the parameter space. However, we can observe how region B (best performance) is extended in Fig. 11 (c). Furthermore, varying $\Delta\phi$ also affects the performance of the tasks at a specific point of $\Delta\omega_{i}/2\pi$ and $\overline{P}_{\textrm{i}}$.

### v2

@@ 6.7.2 Impact of $\Delta\phi$ when addressing multiple, different tasks
When addressing different tasks, changing the value of $\Delta\phi$ does not always have an impact on the performance behavior of each task over the parameter space as shown in Fig. 11 for the NARMA-10 task.
@@
In Fig. 11 (a,b,d), there is not much variation in the performance over the parameter space. However, we can observe how region B (best performance) is extended in Fig. 11 (c). Furthermore, varying $\Delta\phi$ also affects the performance of the tasks at a specific point of $\Delta\omega_{i}/2\pi$ and $\overline{P}_{\textrm{i}}$.
@@
Table 3: Comparison of the best performance of each task for different values of $\Delta\phi$ for a $\overline{P}_{\textrm{T}}$ range of $[-20,25]$ dBm and a $\Delta\omega/2\pi$ range of $[-100,100]$ GHz.